What is the domain and range of the funtion... f(x)=2x+4/(x-7)(x+5)
Answer by lwsshak3(11628)

What is the domain and range of the funtion... f(x)=2x+4/(x-7)(x+5)
***
f(x)=2x+4/(x-7)(x+5)
y=2x+4/(x-7)(x+5)
..
set denominator=0
(x-7)(x+5)=0
x-7=0
x≠7
..
x+5=0
x≠-5
..
x-intercept
set y=0
2x+4=0
x=-2
..
Domain: (-∞, -5) U (-5, 7) U (7, ∞)
Range (-∞, ∞)